Hi, can you confirm my answers, please?

5. Bennings, Inc. paid out total dividends equal to $11,241,200 in 2018. If Jennings reported Retained earnings of $320,868,600 for 2017 and Retained earnings of $350,290,000 for 2018, what was Jennings reported net income in 2018?

Answer: 40,662,600

6. In 2018, Variman, Incorporated had Gross Accounts receivable of $6,456,100 and management estimated the Allowance for Doubtful accounts to be $320,600. Compute the ratio of Allowance of Doubtful accounts to Net Accounts receivable for Variman, Inc. for 2018 (that is, compute Allowance/Net Accounts receivable – record your answer as a percent rounded to one decimal place; for example, record 0.234567 as 23.5%).

Answer: 5.2%

7. In 2018, the Whaddock Company purchased 745,000 units from its supplier at a cost of $160.00 per unit. Whaddock sold 625,000 units of its product in 2018 at a price of $220.00 per unit. Whaddock began 2018 with 125,000 units in inventory (inventory is carried at a cost of $160.00 per unit). Using this information, compute Whaddock’s gross profit for 2018.

Answer: 37,500,000

8. Referring back to the previous question, compute Whaddock’s 2018 ending inventory balance in dollars (valued at cost per unit of $160.00).

Answer: 39,200,000

9. In 2018, Timmers, Inc. (a retail clothing company) sold 57,200 units of its product at an average price of $250.00 per unit. The company reported estimated returns and allowances in 2018 of 1.65 percent of gross revenue. Timmers actually purchased 59,500 units of its product from its manufacturer in 2018 at an average cost of $180.00 per unit. Timmers began 2018 with 8,250 units of its product in inventory (carried at an average cost of $180.00 per unit). Operating expenses (excluding depreciation) for Timmers, Inc. in 2018 were $718,600 and depreciation expense was $310,000. Timmers had $3,000,000 in debt outstanding throughout all of 2018. This debt carried an average interest rate of 5.5 percent. Finally, Timmers’ tax rate was 40 percent. Timmers’ fiscal year runs from January 1 through December 31. Given this information, compute net income for Timmers for 2018.

Answer: 1,544,670

10. Referring back to the previous problem, compute Timmers’ ending inventory balance for 2018 (that is, what did Timmers report as inventory on its December 31, 2018 balance sheet).

Answer: 1,899,000
Answer #1

Question 5)

Dividend payout =$11,241,200

Retain earning in 2018=$350,290,000

Retain earning in 2017=320,868,600

Addition to retained earning in 2018=Retain earning in 2018-Retain earning in 2017=350290000-320868600=$29421400

So net income in 2018=Addition to retained earning in 2018+Dividend payout=29421400+11241200=$40,662,600